OpenStudy (anonymous):

Tan (90+Theta) / 1-Tan 90 Tan Theta

OpenStudy (hba):

\[\frac{ \tan (90+\theta) }{ 1- \tan \theta \tan90 }\]

hartnn (hartnn):

\(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) do you know tan (A+B) formula ?

hartnn (hartnn):

\[\tan(A+B)=\dfrac{\tan A+\tan B }{1-\tan A . \tan B}\] just put A = 90 , B = theta

hartnn (hartnn):

i think the question must be (Tan 90+tan Theta) / (1-Tan 90 Tan Theta)